### v1

# Integrating Window-Based Correlated Decoding with Constant-Time Logical Gates for Large-Scale Quantum Computation

Abstract
Large-scale quantum computation requires to be performed in the fault-tolerant manner. One crucial issue of fault-tolerant quantum computing (FTQC) is reducing the overhead of implementing logical gates. Recently proposed correlated decoding and “algorithmic fault tolerance” achieve fast logical gates that enables universal quantum computation. However, for circuits involving mid-circuit measurements and feedback, this approach is incompatible with window-based decoding, which is a natural requirement for handling large-scale circuits. In this letter, we propose an alternative architecture that employs delayed fixup circuits, integrating window-based correlated decoding with fast transversal gates. This design significantly reduce both the frequency and duration of correlated decoding, while maintaining support for constant-time logical gates and universality across a broad class of quantum codes. More importantly, by spatial parallelism of windows, this architecture well adapts to time-optimal FTQC, making it particularly useful for large-scale computation. Using Shor’s algorithm as an example, we explore the application of our architecture and reveals the promising potential of using fast transversal gates to perform large-scale quantum computing tasks with acceptable overhead on physical systems like ion traps.
Introduction—Quantum computing offers the potential to tackle complex problems such as large integer factorization [1] and quantum simulation [2, 3]. Large-scale quantum computation relies on quantum error correction (QEC) [4, 5] and fault-tolerant quantum computing (FTQC) [6, 7]. In the conventional FTQC framework, error correction follows each logical gate to prevent error accumulation [8], typically involving $\Theta(d)$ rounds of syndrome extraction, where $d$ is the code distance. Consequently, the time of a logical gate implementation is directly relavant to $d$.
Based on correlated decoding, Ref. [9] indicated that only one round of syndrome extraction is required after transversal gates, thereby achieving fast transversal gate. However, transversal gates of any QEC code are not universal [10, 11, 12]. For example, in the surface code, the implementation of logical $T$ gates requires magic states and subsequent fixup operations according to the mid-circuit measurement results [13]. Applying fast transversal gates directly to these circuits can lead to decoding failures due to insufficient syndrome extraction rounds before mid-circuit measurement.
Recently, a fast $T$ gate scheme called “algorithmic fault tolerance” was proposed [14], where certain appropriate logical operators are applied during the decoding to reinterpret inconsistent measurement results of magic states. Finding these logical Pauli operators requires constructing and solving an additional linear system of equations. Particularly, the decoder and feedback process must be invoked after each mid-circuit measurement on magic states (see Fig. 1a). Since correlated decoding algorithms are typically slow, this increases the frequency and duration of decoding.
More notably, this design is not compatible with the window-based decoding [15], which invokes decoder within each fixed-size window. The window-based decoding stands out as a pivotal technology, enabling the decoding of large volumes of syndrome data in a scalable and efficient manner [15, 16, 17, 18, 19]. This approach divides the continuous stream of syndrome information into manageable windows, allowing for parallel processing and reducing the complexity of decoding in large-scale quantum circuits.
In this letter, we design an alternative architecture that realizes constant-time logical gates for universal quantum computing, seamlessly integrating with window-based correlated decoding. This method utilizes delayed fixup circuits to postpone the fixup until sufficient syndrome extraction rounds have occurred, thereby ensuring fault-tolerant measurements and feedback. While achieving fast logical gates and universality, the main advantage of our architecture is that it permits only a single decoder invocation for each sliding window (see Fig. 1b). In other words, syndrome data within a window is processed in a single batch, thus enhancing the efficiency of the correlated decoder in large-scale circuits.
Figure 1: Comparison of correlated decoding without and with window architecture. (a) Without window architecture, correlated decoding is invoked after each mid-circuit measurement. (b) With sliding windows and delayed fixups, correlated decoding processes syndrome data within each fixed-size window.
Despite decoding with temporal sliding-window being well-suited for quantum circuits with sequentially implemented logical gates, time-optimal quantum computation (TOQC), where many logical gates operate in parallel, is another common structure in large-scale quantum computing [20, 21, 22]. TOQC converts time overhead into space overhead, increasing the number of parallel qubits. Since the time for exact correlated decoding grows exponentially with the number of logical qubits [9, 16, 17], it faces significant challenges in handling TOQC circuits with large qubit counts. Interestingly, the spatial window decoding uniquely fits TOQC, which reduces the size of windows and harnesses parallelism to accelerate the decoding, thereby broadens the range of applications for window-based correlated decoding.
Furthermore, we explore the application of our architecture in Shor’s algorithm as an example. Considering connectivity requirements, transversal gates are more feasible on platforms like ion traps [23, 24] or neutral atoms [25]. One QEC round on ion trap platforms is approximately 100 to 1,000 times slower than on superconducting platforms [26, 27, 28], leading to a difference in clock speed of logical gates [29]. For instance, factoring a 2048-bit integer on superconducting systems takes 8 hours [21]. One might infer that ion trap systems could require an impractical time of 800 to 8,000 hours. However, our analysis reveals that the overall runtime of the algorithm in ion-trap systems, when using fast transversal gates and TOQC circuits, is only around an order of magnitude longer than that in superconducting systems, while the space overhead does not increase significantly. Overall, we provide optimistic support for performing large-scale quantum computation tasks with acceptable space-time overhead on physical systems such as ion traps.
Delayed Fixup Circuits—We illustrate the use of delayed fixup circuits for fast implementation of non-Clifford gates, specifically in surface codes, though the approach applies to other QEC codes with transversal CNOT gates and measurements [30, 31]. For surface codes, $T$ gates are implemented using the gadget in Fig. 2a, where fixup operation is classically controlled by the measurement outcome on magic state. When executing the $T$-gadget with fast transversal gates, fewer than $\Theta(d)$ syndrome extraction rounds may be insufficient to ensure fault-tolerant decoding. To accumulate $\Theta(d)$ rounds of syndrome data, computational qubits must repeatedly extract syndromes, followed by decoding and feedback, making $T$ gate implementation time potentially proportional to code distance $d$ in the worst case, rather than constant.
Here, we introduce the delayed fixup circuit to achieve constant-time non-Clifford gates. The delayed fixup circuits like Fig. 2b have been elegantly designed in other previous works [20, 22, 32], whose primary purpose is to achieve gate parallelism for reducing time overhead, rather than to realize constant-time non-Clifford gates. By selecting the ancilla measurement basis, fixup decisions can be delayed. For example, in Fig. 2b, the delayed fixup circuit uses two ancilla states, |$\ket{T}=(\ket{0}+e^{i\frac{\pi}{4}}\ket{1})/{\sqrt{2}}$ and |$\ket{S}=(\ket{0}+i\ket{1})/{\sqrt{2}}$, for implementing the $T$ gate and the $S$ fixup, respectively. On the ancilla state |$\ket{S}$, the measurement is delayed until sufficient syndrome extraction rounds have been completed. Once correlated decoding provides a reliable measurement result for |$\ket{T}$, the basis of |$\ket{S}$ measurement is chosen. For circuits with multiple non-commuting $T$ gates, the |$\ket{S}$ measurement basis must be chosen sequentially. Note that on the computational qubits, only transversal CNOT gates are performed, which eliminates repeated syndrome extraction and ensures that the implementation time of the $T$ gate is irrelevant to $d$.
Figure 2: (a) Gadget for implementing $T$ gate. (b) $T$-gadget with delayed $S$ fixup. The decision for $S$ fixup is delayed until the choice of the $X$ or $Z$ measurement basis.
In summary, the delayed fixup circuit ensures sufficient syndrome extraction rounds before fixup operations, making the decision on fixup fault-tolerant. Combined with transversal Clifford gates on surface codes, it is easy to confirm that this architecture achieves constant-time logical gates in a universal set. A potential concern is that keeping the |$\ket{S}$ state longer in the circuit might accumulate logical errors or increase qubit usage. However, as we will show later, with an appropriate window length, the |$\ket{S}$ state persists for at most $\Theta(d)$ rounds, preventing significant increase in logical error rates or resource overhead.
Correlated Decoding with Sliding Windows—Next, we introduce the temporal sliding windows in correlated decoding and explain its unique adaptability for delayed fixup circuits. For large-scale quantum circuits, dividing syndrome data into windows for separate processing is essential. However, applying a sliding-window architecture to fast non-Clifford gate circuits is not straightforward. The sliding window method processes entire syndrome data in each window, which is clearly incompatible with the fast $T$-gate scheme in Ref. [14], where correlated decoding is applied after each magic state measurement.
In contrast, the sliding windows works well in delayed fixup circuits. In a feed-forward sliding window architecture, the window with length $n_{w}$ slides forward in time (see Fig. 3a). At each step, correlated decoding is performed on $n_{w}$ rounds of syndrome data. The decoding result of each window is divided into two regions: a commit region and a buffer region with lengths $n_{c}$ and $n_{b}$, respectively, where $n_{w}=n_{c}+n_{b}$. After that, the window slides forward by $n_{c}$ rounds. In other words, the buffer region is the overlap between two adjacent windows. To ensure fault tolerance of the decoding, $n_{w}$, $n_{c}$, and $n_{b}$ are typically proportional to $\Theta(d)$.
Since the syndrome checks near the window boundary lack future syndrome information, they might be matched incorrectly. Thus, only the decoding results within the commit region, which is sufficiently distant from the boundary, are considered reliable. Specifically, if all vertices of a hyperedge lie entirely within the commit region or spanning two regions, the result is preserved as final and the syndrome checks on the vertices are flipped. If all vertices of a hyperedge are entirely within the buffer region, it is deferred for the next window to resolve.
Figure 3: Sliding window architecture and application in $T$-gadget. (a) Temporally feed-forward sliding window. Syndromes and hyperedges from each correlated decoding are represented within a window. The windows are processed sequentially, with each step removing the green hyperedge. (b) The measurement of |$\ket{T}$ occurs in the commit region of the second window, followed by the immediate measurement of |$\ket{S}$. In the third window, |$\ket{S}$ is independently decoded, with the state persisting for at most one window length.
Then, we can discuss circuits containing mid-circuit measurements and feedback more concretely. Taking the $T$-gadget in Fig. 3b as an example, when state |$\ket{T}$ measurement appears in the commit region of a sliding window, the decoding yields a reliable measurement result. Following this, the state |$\ket{S}$ is measured in the appropriate basis immediately. The measurement outcome of |$\ket{S}$ will depend on the decoding result from the subsequent window. In particular, in the subsequent window, since the state |$\ket{S}$ does not interact with other qubits via two-qubit gates, it can be decoded independently as a single logical qubit. Note that independent decoding of |$\ket{S}$ states is crucial, especially when multiple non-commuting $T$ gates with $S$ fixup are performed consecutively. If the previous |$\ket{S}$ state is decoded within an entire window, the next |$\ket{S}$ measurement must be delayed until the following window. For a sequence of $S$ fixups, this will clearly result in an accumulation of errors.
This example also highlights an additional advantage of window-based correlated decoding. When there are no two-qubit gate interactions between subsets of qubits within a window, the correlated decoding can be partitioned into several independent parts. This partitioning decreases the number of variables in the correlated decoding and enables further time reductions through parallelization. Essentially, this represents a specific instance of the spatial parallel window decoding, which will be discussed later.
Spatial Parallel Windows—Our starting point is the concept of TOQC for large-scale quantum computing, first proposed in Ref. [20]. Fig. 4a illustrates an circuit of TOQC, containing teleportation and delayed fixup circuits. The implementation time of TOQC is primarily dominated by the depth of measurements and classical feedback, thus reducing time overhead in large-scale computing tasks.
Figure 4: Application of Spatial Windows in TOQC. (a) Parallel Clifford+$T$ gates in TOQC. The blue boxes divide the space into multiple blocks. The state |$\ket{S}$ in the $T$ gadget and its delayed measurement are also highlighted. (b) When the temporal window is shorter than the spacing between transversal CNOT gates, the window divides into independent parts naturally. (c) If the temporal window covers the entire circuit, it can be spatially divided, with syndrome data processed in a feed-forward manner. (d) In parallel spatial windows, decoding occurs in two stages, retaining green hyperedges as results in each.
Although TOQC reduces the time overhead, it does not lower the overall space-time cost. For instance, executing $\mathcal{O}(k)$ groups of logical gates in parallel increases the number of qubits by $\mathcal{O}(k)$ times. This naturally leads to challenges in correlated decoding of TOQC circuits. As discussed before, for circuits with large depth, correlated decoding can avoid processing too much syndrome data at once by using temporal sliding windows. However, in TOQC circuits, the circuit depth is converted into a larger number of qubits. Given that the complexity of exact correlated decoding increases exponentially with the number of qubits, directly applying correlated decoding to large TOQC circuits presents significant limitations.
Here, similar to temporal windows, we find that TOQC circuits can be divided into spatial windows for correlated decoding. Notice that TOQC circuits can naturally be divided into blocks spatially, each comprising Bell state preparations, logical Clifford+$T$ gates, and Bell-basis measurements and feedback (see Fig. 4a). The blocks are interconnected by the initial and final transversal CNOT gates.
Now let us explain spatial windows in two scenarios (see Fig. 4b and 4c). First, if the depth $L$ of logical Clifford+$T$ gates in each block satisfies $L>n_{w}$ (recall that $n_{w}$ is the length of the temporal window), the initial and final transversal CNOT gates will not be in the same window. Consequently, this allows the temporal window to be divided into several independent parts for parallel correlated decoding. Conversely, if $L\leq n_{w}$, the temporal window cannot be directly divided into independent parts. In this case, we further divide the windows spatially, with each window containing syndrome data from $m_{w}$ blocks. Similarly, each window is composed of a commit region and a buffer region, containing $m_{c}$ and $m_{b}$ blocks, respectively. The intuition for spatial windows is that syndrome information between blocks that are spatially far apart becomes nearly independent. To keep the commit region away from the window boundary, the size of the buffer region is set to satisfy $m_{b}L=\Theta(d)$. In this setting, an error string connecting the commit region to the window boundary will contain at least $m_{b}L$ hyperedges. The condition $m_{b}L=\Theta(d)$ ensures such errors occur with exponentially low probability as $d$ increases and also confirms the comparable sizes of the temporal and spatial windows.
Spatial window decoding can be executed in a feed-forward manner. First, correlated decoding is applied to the syndrome data within the window, preserving hyperedges located entirely within the commit region or those spanning both the commit and buffer regions. Subsequently, the syndrome checks on the vertices of these hyperedges are flipped.
Furthermore, spatial windows can be constructed and processed in parallel to reduce decoding time (see Fig. 4d). Each window is divided into three parts: a central commit region, flanked by buffer regions on both sides. Similarly, the buffer regions are sized to satisfy $m_{b}L=\Theta(d)$. Correlated decoding is performed once for each window and then the processed commit regions naturally divide the remaining areas into several independent parts. Consequently, once more decoding can be performed in parallel for each part. This process is similar to the parallel window method in Ref. [18], except that the windows here are divided spatially.
Numerical Results—We present two examples (see details in Supplementary Material) to illustrate the performance of the window-based correlated decoding. In the first example, the length of temporal window is $n_{c}=n_{b}=\frac{d+1}{2}$. In the second, we test both feed-forward and parallel spatial windows with $m_{c}=1$ in both cases. Our setup ensures that the length of error strings from the commit region to the window boundary is at least $\frac{d+1}{2}$. Compared to correlated decoding on the syndrome data of the entire circuit, the performance of window-based correlated decoding is almost identical (see Fig. 5).
Figure 5: Decoding failure probabilities $p_{\text{fail}}$ with and without window architecture. The decoding methods are distinguished by colors, and the different code distances are represented by varying shades. (a) Comparison of $p_{\text{fail}}$ with and without temporal feed-forward window decoding. (b) Comparison of $p_{\text{fail}}$ with and without spatial window decoding, including both feed-forward and parallel types.
Large-Scale Quantum Computation—Let us consider applying our architecture to large-scale quantum computation, using Shor’s algorithm as an example. Our analysis builds upon Ref. [21], comparing the time and space overhead of implementing Shor’s algorithm using fast transversal gates in ion trap systems versus the lattice surgery [33, 34, 22] in superconducting systems.
Some basic assumptions include the following. First, it is assumed that the error rates for physical gates, initialization and measurements are comparable between the two systems, while the ion-trap system can provide required connectivity. Second, the QEC round time for the ion-trap system is 100 to 1,000 times that of the superconducting system [26, 27, 28]. Additionally, in ion trap systems, the reaction time, defined by measurement plus classical processing and feedback time, is about an order of magnitude longer than in superconducting systems due to longer measurement time [32, 35, 36, 37].
Ref. [21] identifies the two most time-consuming subroutines as the ripple-carry adder [38] and the lookup table [39]. The ripple-carry adder, executable via TOQC circuits, is limited by the reaction time. At this point, fast transversal gates offer negligible time savings since the computation time is proportional to the measurement depth. In contrast, the implementation time of lookup table is dominated by multi-target CNOT gates, also known as Clifford-limited [32]. With fast transversal gates and ancilla cat states, multi-target CNOT gates can be implemented in a single QEC round, mitigating the effect of slower QEC rounds in ion trap systems by approximately a factor of $d$. Given the above factors, the total time of these two subroutines is approximately an order of magnitude longer in ion trap systems (see Supplementary Material for a more specific calculation).
Moreover, fast transversal gates also reduce the space overhead. Since the algorithm is executed with fewer QEC rounds by fast transversal gates, the accumulated logical error rate on the computational qubits will be lower, thus relaxing the requirement for the code distance $d$. This reduction in space overhead is quite significant, approximately 14% when reducing the code distance $d$ from 27 to 25 (see Supplementary Material). Additionally, systems with flexible connectivity will further reduce qubit requirements by saving routing qubits and allowing flexible qubit reuse.
However, the space occupied by magic state distillation must be considered [40]. The differing ratios of QEC round time and reaction time between ion trap and superconducting systems lead to variations in the generation and consumption rates of magic states. In ion trap system with fast transversal gates and TOQC circuits, magic states will be consumed faster, measured in units of QEC rounds, which indicates the distillation rate needs to increase. Fortunately, fast transversal gates applied in distillation can theoretically reduce space-time overhead by $\Theta(d)$ times [14]. Given that distillation protocol based on lattice surgery is highly optimized [41], the actual factor of reduction in space-time overhead will possibly be less than $d$. Considering these factors comprehensively, although distillation factories require more space, the overhead will remain acceptable. Overall, using fast transversal gates in ion trap systems yields space overhead of Shor’s algorithm comparable to using lattice surgery in superconducting systems [21], as corroborated by the calculations in the Supplementary Material.
Discussion—We present an architecture that integrates window-based correlated decoding with fast logical gates for universal quantum computing, offering broad applicability in large-scale quantum computing scenarios. The adaptation to temporal windows for parallelism, as seen in Refs. [18, 19], is straightforward. Nevertheless, investigating the general connection between spatial windows parallelism and the circuit structures remains an interesting question. Additionally, preliminary analysis of Shor’s algorithm in ion trap systems suggests promising potential, with more detailed resource evaluations and optimizations to be explored in future work.
Acknowledgements.
We thank the Gurobi Optimization team for providing the Gurobi software package [42], which was used for solving optimization problems in this work. This work was supported by the National Natural Science Foundation of China (Grant No. 12034018).
Supplemental Material
I correlated decoding within a window
In this section, we provide a explicit description of correlated decoding within a window. For CSS codes [6], the logical space is determined by stabilizer generators, denoted as $s_{x}^{(i)}$ and $s_{z}^{(i)}$, which represent $i$-th $X$-type and $Z$-type stabilizer generators, respectively. When running multiple rounds of syndrome extraction circuits, the error syndrome $\sigma_{n,k}(s_{x/z}^{(i)})\in\{+1,-1\}$ is defined as the measurement result of the stabilizer generator on the $n$-th logical qubit in the $k$-th round.
Next, the syndrome checks (or checks in brief) $C_{n,k}(s_{x/z}^{(i)})$, with value $c_{n,k}(s_{x/z}^{(i)})\in\{+1,-1\}$, are defined as the product of $\sigma_{n,k}(s_{x/z}^{(i)})$ and its previous result (if one exists), obtained by back-propagating the stabilizer operator through the circuit to the point where it was previously measured. Specifically, after the $k$-th round of syndrome extraction, if logical $I_{L}$, $X_{L}$, $Y_{L}$, or $Z_{L}$ gate is applied, then
$c_{n,k}(s_{x/z}^{(i)})=\sigma_{n,k}(s_{x/z}^{(i)})\sigma_{n,k-1}(s_{x/z}^{(i)}).$ (S1)
If logical $H_{L}$ gate is applied after the $k$-th round of syndrome extraction, then
$c_{n,k}(s_{x/z}^{(i)})=\sigma_{n,k}(s_{x/z}^{(i)})\sigma_{n,k-1}(s_{z/x}^{(i)}).$ (S2)
If logical CNOT gate between qubit $n_{1}$ and $n_{2}$ is applied after the $k$-th round of syndrome extraction, then
$c_{n_{1},k}(s_{x}^{(i)})$ $=\sigma_{n_{1},k}(s_{x}^{(i)})\sigma_{n_{1},k-1}(s_{x}^{(i)})\sigma_{n_{2},k-1}(s_{x}^{(i)}),$ (S3)
$c_{n_{2},k}(s_{x}^{(i)})$ $=\sigma_{n_{2},k}(s_{x}^{(i)})\sigma_{n_{2},k-1}(s_{x}^{(i)}),$
$c_{n_{1},k}(s_{z}^{(i)})$ $=\sigma_{n_{1},k}(s_{z}^{(i)})\sigma_{n_{1},k-1}(s_{z}^{(i)}),$
$c_{n_{2},k}(s_{z}^{(i)})$ $=\sigma_{n_{2},k}(s_{z}^{(i)})\sigma_{n_{2},k-1}(s_{z}^{(i)})\sigma_{n_{1},k-1}(s_{z}^{(i)}).$
In correlated decoding, each error event $E_{j}$ in the circuit is assigned an error probability $p_{j}$ and a set of checks $C_{j}=\{C_{n,k}(s_{x/z}^{(i)})|c_{n,k}(s_{x/z}^{(i)})=-1\text{ if }E_{j}\text{occurs}\}$. Each error event $E_{j}$ naturally corresponds to a hyperedge $e_{j}$ with a weight of $-\log p_{j}$, whose vertices correspond to the checks in $C_{j}$. Consequently, the correlated decoding problem is transformed into a minimum matching problem on a weighted hypergraph. According to Ref. [9], the solution to this problem can be exactly solved by a mixed-integer programming algorithm, known as the most-likely error (MLE) decoder, or approximately solved by hypergraph union-find (HUF) decoder [43, 44].
Next, let us describe how correlated decoding is executed within a window. First, we need to establish the relationship between each error event $E_{j}$ with its error rate $p_{j}$ and its corresponding hyperedge $e_{j}$. For each error event $E_{j}$, a noiseless circuit with the error event $E_{j}$ is simulated, and the set of checks with a value of -1, denoted as $C_{j}$, is recorded. However, in a window of limited size, some checks in $C_{j}$ may fall outside the window. Thus, in correlated decoding within a window, we remove checks that lie outside the current window and consider the remaining checks as the vertices of hyperedge $e_{j}$. If none of the vertices are within the current window, the error event $E_{j}$ is removed from the error event set of the current window. Additionally, hyperedges spanning both the commit and buffer regions will be removed after decoding. Therefore, for hyperedges spanning the commit and buffer regions of a window, we only consider the hyperedge $e_{j}$ and the corresponding error event $E_{j}$ in the current window, and remove $E_{j}$ from the error event set of the next window.
After each correlated decoding within a window, we preserve the hyperedges inside the commit region and those on the boundary of the commit and buffer region as reliable results. The syndrome information corresponding to these hyperedges is then eliminated, meaning the check values at the vertices of these hyperedges are flipped. The decoding of subsequent windows is then carried out, and the final output consists of all reliable hyperedges in each decoding as the complete result.
II size of temporal and spatial windows
For spatial windows, their size is immediately apparent. Each spatial window processes syndrome data from $\Theta(d)$ rounds on $n$ qubits at a time, so the size of correlated decoding remains acceptable. However, in TOQC circuits, a spatial window includes multiple $n$-qubit blocks. One might wonder whether the size of a spatial window remains acceptable. In the main text, we pointed out that the sizes of temporal and spatial windows are comparable. Here, we will more clearly demonstrate this conclusion.
To ensure fault tolerance, the distance between the commit region and the window boundary must be sufficiently large. This distance refers to the minimum number of hyperedges required to connect two vertices. In fact, this distance only depends on the size of the buffer region. For temporal and spatial windows, $n_{b}$ and $m_{b}$ can be set to $(d+1)/2$ and $(d+1)/2L$, respectively. This ensures that the probability of an error chain connecting the commit region to the window boundary is $\mathcal{O}(p^{(d+1)/2})$, which will not significantly contribute to the logical error rate $P_{L}$. On the other hand, the size of the commit region is not restricted. Typically, to output as many reliable decoding results as possible while keeping the overall window size manageable, the size of the commit region is set equal to that of the buffer region.
Now, under these settings, let us calculate the sizes of the temporal and spatial windows. Suppose the temporal window includes $n$ qubits in the circuit, then the space-time volume of the temporal window is $n(n_{c}+n_{b})=n(d+1)$. For the spatial window, assume that each block in the TOQC circuit has $n$ qubits and a logical gate depth of $L$. The space-time volume of the spatial window is also $n(m_{c}+m_{b})L=n(d+1)$. This result dispels concerns about the size of spatial windows, demonstrating that both spatial and temporal windows can perform correlated decoding within a limited scale.
III Time-optimal quantum computation (TOQC) circuit
Here, we explain additional details of the TOQC circuit mentioned in the main text. Generally, a TOQC circuit consists of two components, the teleportation circuit (Fig. S1) and the delayed choice circuit. When the input state $\psi$ is an $n$-qubit state, the operation $U_{i}$ is applied to a block composed of $n$ Bell pairs. Afterward, Bell basis measurements are performed, and based on the theory of teleportation, it can be easily verified that $U_{i}$ is successfully applied to the state $\psi$, up to a recovery operation $R$. When $U_{i}$ is a Clifford operation, the recovery operator $R$ is an element of the n-qubit Pauli group.
When $U_{i}$ is a Clifford operation combined with a $T$-gadget, delayed $S$ fixup can be implement to ensure that the recovery operator $R$ remains a Pauli operator and can be tracked in the Pauli frame[45, 46]. The recovery operator $R$ will affect subsequent recovery operators and the measurement basis of the $S$ fixups. For example, if the recovery operator $X$ is generated, the equation $XTX=T^{\dagger}=TSZ$ implies that the $S$ gate and $Z$ gate must be applied afterward, which will be reflected in the measurement basis for future $S$ fixups and in the Pauli frame. Therefore, at the end of the circuit, the measurement basis of the $S$ state will be determined sequentially.
In the Shor’s algorithm example mentioned in the main text, the non-Clifford gates in the TOQC circuit are typically Toffoli gates. The parallel execution of Toffoli gates can be achieved using delayed CZ gates with the similar methods in the TOQC, as discussed later.
Additionally, note that the TOQC circuit discussed in the main text is not entirely identical to the results in Ref. [20], which Ref. [20] can be regarded as a more general form of TOQC. In the general form, spatial window decoding is also applicable. This is because the general TOQC structure shares similarities with the one discussed in the main text. Specifically, the TOQC circuit can be divided into spatial blocks, and transversal CNOT gates between blocks only appear at the beginning or end of the circuit. As a result, the discussion in the main text is also valid for the TOQC described in Ref. [20].
Figure S1: Teleportation circuit in TOQC.
IV More details of numerical simulations
Figure S2: Details of the two examples in the numerical simulation. (a) Circuit and feed-forward sliding window for the first example. (b) Circuit for the second example. (c) Feed-forward spatial windows. (d) Parallel spatial windows.
Numerical simulations are conducted on two examples of sliding window applications in the main text. Here, we provide a clearer description of their details in Fig. S2. The logical qubits in the circuit are encoded on surface codes, enabling transversal CNOT gates and single-qubit Clifford gates from the set $\{X,Y,Z,H\}$. Fig. S2 a-b illustrate the specific circuits for both examples, where syndrome extraction is performed for single round after each logical gate. The division of windows is shown in Fig. S2 a and c-d.
In the first example, the circuit consists of two logical qubits and a series of transversal CNOT gates and single-qubit Clifford layers, with a total depth of $23(d+1)/$ 2. The sliding window length is set to $n_{w}=d+1$, where the commit and buffer regions each occupy half. The entire circuit is decoded within two temporal windows sequentially.
In the second example, the circuit consists of five logical qubits and has a structure similar to TOQC with $L=(d+1)/2$. We tested the decoding process under two window architectures. In the feed-forward window architecture, each window contains syndrome data from two blocks. Correlated decoding is performed sequentially across four windows, with the window sliding by one block in the spatial direction each time. In the parallel window architecture, decoding is carried out in two steps: first, two disjoint spatial windows are decoded, followed by independent correlated decoding in the remaining buffer regions of the two parts.
We simulate the circuits under circuit-level noise model to obtained syndrome data. Specifically, the depolarizing Pauli noise channels are defined as follows:
$\mathcal{E}_{1}(\rho_{1})$ $=(1-p)\rho_{1}+(p/3)\sum_{P\in\{X,Y,Z\}}P\rho_{1}P,$ (S4)
$\mathcal{E}_{2}(\rho_{2})$ $=(1-p)\rho_{2}+(p/15)\times$
$\sum_{\begin{subarray}{c}P_{1},P_{2}\in\{I,X,Y,Z\},\\ P_{1}\otimes P_{2}\neq I\otimes I\end{subarray}}P_{1}\otimes P_{2}\rho_{2}P_{1}\otimes P_{2},$
where $\rho_{1}$ and $\rho_{2}$ are single-qubit and two-qubit density matrices respectively. In the simulated circuits, we apply $\mathcal{E}_{1}$ after single-qubit gates, and idle operations, and $\mathcal{E}_{2}$ after two-qubit gates. Additionally, the measurement result and state initialization flips with the probability of $p$. We simulate the circuits in the Heisenberg representation [47], randomly inserting Pauli noise with corresponding probability. These Clifford circuits can be efficiently simulated in polynomial time, according to the Gottesman-Knill theorem [8, 48].
Then, correlated decoding is applied within each window. The decoding process uses an MLE decoder to obtain the exact solution for correlated decoding, implemented through the Gurobi solver [42]. One difference from the original correlated decoding literature is that we pre-merge several error events $\{E_{i}\}$ corresponding to the same hyperedge and assign an error probability based on the likelihood of an odd number of these events occurring. This approach accounts for the equivalence of error events. We expect the results after this pre-merging process to be closer to the exact solution for the decoding problem, i.e., most-likelihood decoding [49]. Additionally, the reduction in variables within the ${E_{i}}$ set implies a lower decoding time complexity.
[…]
V implementation of ripple-carry adder and lookup table
In the main text, we discussed the two most time-consuming subroutines in Shor’s algorithm: the ripple-carry adder [38] and the lookup table (or QROM) [39]. Their specific circuit implementations can be found in Ref. [32] and Ref. [50]. The non-Clifford gate in the circuits of these two subroutines is the Toffoli gate. For surface codes and other quantum error correction codes with transversal CNOT gates, the Toffoli gate can be implemented using the gadget circuit in Fig. S3a, where the fixups are CZ and CNOT gates. To construct a TOQC circuit containing Toffoli gates, one can use the delayed CZ correction circuit shown in Fig. S3b, or equivalently implement a CCZ gate using the auto-CZZ ancilla state from Ref. [32].
[…]
For the lookup table, another complexity lies in the implementation of multi-target CNOT gates. Without connectivity constraints, multi-target CNOT gates can be implemented using ancilla cat state (|$(\ket{0}^{\otimes n}+\ket{1}^{\otimes n})/\sqrt{2^{n}}$ followed by two layers of CNOT gates (only one layer on the computaional qubit), as shown in Fig. S4a. Additionally, the $n$-qubit cat state can be prepared in $\lceil\log n\rceil$ layers of CNOT gates, as shown in Fig. S4b. Note that in solid-state quantum systems, such as superconducting system, the time complexity of preparing an $n$-qubit cat state is $O(n)$, assuming only nearest-neighbor physical CNOT gates are allowed. Furthermore, multi-target CNOT gates cannot be heavily parallelized, as routing qubits will be occupied.
Figure S4: (a) Implementation of multi-target CNOT gate with ancilla cat state. (b) Circuit for preparing $n$-qubit cat state in $\lceil\log n\rceil$ layers.
VI A More Specific Estimate of the Space-Time Overhead for Shor’s Algorithm
Our estimate follows the framework laid out in Ref. [21], which elaborates on the execution process of the Shor’s algorithm within the lattice surgery framework of superconducting systems. Actually, we did not conduct a comprehensive analysis of all detailed steps of Shor’s algorithm, but instead focused on the most time-consuming part, the lookup addition, which accounts for the vast majority of the space-time overhead when factoring large integers.
Our evaluation is not a complete reconstruction but rather focuses on comparing with the lattice surgery method, examining how fast transversal gates impact the overhead in various aspects. At the algorithm level, the analysis is the same as Ref. [21]. In other words, we maintain the same process flow of the algorithm as Ref. [21] but employ fast transversal gates rather than lattice surgery for the circuit implementation. In Ref. [21], several optimizations are made for the practical execution of Shor’s algorithm to factor a 2048-bit integer. Note that the parameters in this process are selected based on lattice surgery techniques for only superconducting systems. Thus, our analysis serves as a rough reference, leaving room for potential further optimization.
First, we clarify several important assumptions. It is assumed that ion trap systems allow arbitrary connectivity while maintaining the same error rate as superconducting systems. Besides, the time required for one round of QEC in ion trap systems is assumed to be $\gamma$ times longer than that of superconducting systems. Based on existing experimental data and theoretical projections, a $\gamma$ value between $10^{2}$ and $10^{3}$ is a reasonable assumption [26, 27, 28]. Moreover, the reaction time in ion trap systems is estimated to be $\eta$ times that of superconducting systems. In superconducting systems, the reaction time is mainly constrained by the time needed for classical processing and feedback, generally estimated to be around 10 $\mu$[32, 21]. For ion trap systems, the reaction time encompasses both the measurement and classical processing, and feedback, estimated to range between 20 and 200 $\mu$[35, 36, 37]. Therefore, $\eta$ is assumed to fall between 2 and 20. For simplicity, we assume that $\gamma$ and $\eta$ scale proportionally, i.e., $\gamma/\eta=50$. Additionally, the physical error rate is assumed to be $10^{-3}$, and the code distance for the computational qubits is set at $d=27$, which, as shown in Ref. [21], can provide a sufficiently low logical error rate.
Now, we can address the time overhead of Shor’s algorithm. The time complexity is mainly driven by numerous lookup addition operations, including both the lookup phase and the addition phase. According to Ref. [21] and Ref. [32], the time for the lookup phase is roughly proportional to the time taken for multi-target CNOT gates, estimated at 14 ms in superconducting systems. As discussed earlier, with fast transversal gates, multi-target CNOT gates can be executed in a single QEC round (rather than $d$ rounds) on the computational qubit. Thus, the lookup phase time in ion trap systems is estimated to be $14\text{ms}\times\gamma/d=$ 51.9 to 519 ms. In the addition phase, since the circuit is implemented using TOQC, the reduction in time overhead from fast transversal gates is almost negligible. Therefore, the addition phase time in ion trap systems is estimated to be $22\text{ms}\times\eta=$ 44 to 440 ms, where 22 ms is the corresponding time for superconducting systems.
Summing these two phases, the total time for the lookup addition operation in ion trap systems is 2.7 to 27 times longer than in superconducting systems. This ratio can serve as a rough estimate of the overall time difference between the two systems when factoring a large integer using Shor’s algorithm, with ion trap systems requiring approximately one order of magnitude more time.
Regarding space overhead, we mainly focus on optimizing the computational qubits and the preparation of auto-CCZ states. In other words, we assume that the space overhead for other parts (such as registers in the lookup table) remains comparable. This assumption overlooks the differences in routing qubits between the two systems, making it more conservative for ion trap systems.
First, according to the setup in Ref. [21], factoring a 2048-bit integer requires 6,200 computational logical qubits. As mentioned in the main text, the advantage of fast transversal gates lies in their ability to reduce the error rate. For surface codes, the logical error rate per QEC round is estimated as $P_{L}(d)\propto(p/p_{th})^{(d+1)/2}$ [13], where we assume the physical error rate $p$ and the threshold $p_{th}$ are $10^{-3}$ and $10^{-2}$, respectively. We have calculated that, in ion trap systems, a lookup addition takes 95.9-959 ms, which includes 959 QEC rounds. In contrast, this value is $36\text{ms}/1\mu\text{s}$= 36,000 rounds in superconducting systems. This implies a reduction in the cumulative logical error rate of computational qubits by approximately 37 times. Consequently, the code distance can be reduced from 27 to 25, resulting in about a 14% reduction in space overhead, while the overall error rate remains no higher than that in superconducting systems. One might worry that Toffoli gates are excluded from this discussion, as their error rate is also affected by the fidelity of CCZ states. However, note that when factoring large integers using Shor’s algorithm, CCZ states typically require two levels of distillation [21]. After two levels of distillation, the fidelity of the output CCZ states is dominated by the Clifford operations, rather than the fidelity of the input states in the distillation protocol [40, 41]. Therefore, by implementing these Clifford operations with fast transversal gates, the equation of error rate still applies to Toffoli gates.
Additionally, the preparation of auto-CCZ states also contributes to space overhead. Note that these qubits can be reused. Therefore, assuming the required time for this part remains constant, reducing space-time overhead directly implies a reduction in space overhead. First, high-fidelity CCZ states are obtained through two levels of distillation [40]. In superconducting systems, this roughly requires $16\times 113$ logical qubits. Compared to the original Ref. [40], this overhead has already been reduced by about 8 times in Ref. [41]. With fast transversal gates, this overhead is expected to be further reduced by a factor of $\Theta(d)$ [14], which we conservatively estimate as 5 times. Second, the CCZ states need to be converted into auto-CCZ states, a process that involves several CZ gates [32]. In superconducting systems, this process requires around $6\times 113$ logical qubits. Fast transversal gates can speed up these CZ gates by a factor of $d$, which is equivalent to reducing space overhead by a factor of $d$. Finally, the space overhead for preparing auto-CCZ states needs to be multiplied by $\gamma/\eta$. This is because the speed of generating and consuming auto-CCZ states is proportional to $\gamma$ and $\eta$ (at most), respectively. Thus, space overhead needs to be expanded by $\gamma/\eta$ to match the generation and consumption rates of auto-CCZ states. After all calculations, the space overhead in ion trap systems is actually lower, roughly 99.6% of that in superconducting systems.
Note that these results are based on several conservative assumptions. Clearly, these evaluations are quite rough and could be further refined. However, as we have pointed out, a particularly thorough analysis is beyond the scope of this work and will be left for future studies.